OpenStudy (anonymous):

What is the axis of symmetry for the function? y = x2 + 3x - 4 y = -2 y = -1.5 x = -2 x = -1.5

jimthompson5910 (jim_thompson5910):

use the same formula from the last post

jimthompson5910 (jim_thompson5910):

tell me what you get

jimthompson5910 (jim_thompson5910):

if you need it again, it is x = -b/(2a)

OpenStudy (anonymous):

okay just a sec.

jimthompson5910 (jim_thompson5910):

alright

OpenStudy (anonymous):

is a=2 or x

jimthompson5910 (jim_thompson5910):

x^2 + 3x - 4 is the same as 1x^2 + 3x - 4

jimthompson5910 (jim_thompson5910):

so in this case a = 1, b = 3, c = -4

OpenStudy (anonymous):

x=1.5

jimthompson5910 (jim_thompson5910):

x = -b/(2a) x = -3/(2*1) ... plug in a = 1 and b = 3 x = -3/2 x = -1.5 so you are off by a sign (lost the negative sign somewhere)

OpenStudy (anonymous):

so it is -1.5, i thought i was suppose to change it to a possative?

jimthompson5910 (jim_thompson5910):

no the axis of symmetry can be negative

jimthompson5910 (jim_thompson5910):

it's not always positive

jimthompson5910 (jim_thompson5910):

you might be thinking of absolute value?

OpenStudy (anonymous):

o yea i am sorry okay thank you, i might need help later so look for my questions please:)

jimthompson5910 (jim_thompson5910):

sure thing, glad to be of help

jimthompson5910 (jim_thompson5910):

and don't be sorry, silly mistake